# Sailor: Automating Distributed Training over Dynamic, Heterogeneous, and Geo-distributed Clusters

### v1

@@ 3.2. Challenges with Heterogeneous ML Clusters
C1: Quickly searching a vast configuration space. Considering heterogeneous and geo-distributed GPUs creates a vast and complex search space. The ML developer needs to decide how many GPUs to use and how to group them across VMs. This complexity increases further, when accounting for job parallelization plans within each allocation. Furthermore, the optimal allocation and partitioning strategies depend on user objectives and constraints. A planner needs to quickly navigate the large configuration space to adapt to dynamic resource availability in cloud and on-premise environments (Thorpe23Bamboo, ; guo2024cephalo, ; weng22mlaas, ; Gu19Tiresias, ), since maximizing throughput requires adapting parallelization strategies with changes to cluster topology (Athlur22Varuna, ; wagenlander24tenplex, ).
Table 1 shows that Metis (Um24Metis, ), FlashFlex (yan2024flashflex, ), Cephalo (guo2024cephalo, ), Atlas (palak2024atlas, ), and DTFM (yuan2023decentralized, ) explore parts of this large search space. Atlas (palak2024atlas, ) and DTFM (yuan2023decentralized, ) target geo-distributed training, but do not consider heterogeneous GPU types, and do not decide the various parallelism degrees: instead, they take as input the parallelism degrees, and assign these degrees in the available zones. On the other hand, Metis (Um24Metis, ), FlashFlex (yan2024flashflex, ) and Cephalo (guo2024cephalo, ), consider heterogeneous GPU types, but overlook geo-distributed training, and are quite inefficient for dynamic environments. Metis needs a few hours to devise a plan for a 16-GPU cluster (A100 and V100)222with max_permutation_length and device group variance set to 10 and 0.5 respectively, according to the paper (Um24Metis, ), making frequent reevaluation infeasible as GPU availability changes. Cephalo (guo2024cephalo, ) has a search time of  300 sec on a cluster of 64 GPUs333reported on the paper (guo2024cephalo, ), but it is limited only to Fully Sharded Data Parallelism. FlashFlex (yan2024flashflex, ) has a short runtime, but provides inaccurate runtime estimations, leading to suboptimal plans. Furthermore, these planners only optimize for throughput, ignoring budget constraints, and cost (dollars per iteration), which affect the optimal configuration (§5.2.4).
@@
Figure 3. Peak Memory estimations of various baselines compared to the actual peak memory, for the OPT-350M model on a homogeneous cluster of 4 Grace-Hopper per node. $N$ stands for the number of nodes, $gbs$ is the global batch size, $pp$ pipeline parallelism, $tp$ tensor parallelism, $dp$ data parallelism, and $mbs$ the microbatch size.
@@
Unfortunately, these estimations are often inaccurate, resulting in suboptimal or invalid plans. First, planners either completely ignore memory footprint (li2022amp, ), or underestimate the amount of memory requirements during training (Athlur22Varuna, ; Jang2023Oobleck, ), omitting activations, optimizer states, and memory fragmentation, or assume the training memory footprint is uniform across all devices and pipeline stages (yan2024flashflex, ; Lin24nnScaler, ). As a result, these systems may find plans that cause OOM errors, when deployed. Figure 3 compares memory footprint estimations with the real footprint on a homogeneous cluster of up to 16 Grace-Hopper nodes for the OPT-350M model, showing that planners can be 25-95% off when estimating memory footprint. Second, planners often poorly model training time, due to incorrect assumptions about network bandwidth and ignoring communication-computation overlap (Zhang17Poseidon, ; megatron-deepspeed, ).
@@
Modeling memory footprint and iteration time becomes even more complex with heterogeneity (Mo24Heet, ). Different GPU generations vary in compute performance (e.g., TFLOPs), and memory capacity (zhu2024nanoflow, ). Thus, a configuration that fits in one GPU, might cause OOM errors in another GPU. Additionally, stragglers and network bandwidth differences (especially in geo-distributed setups (strati24crossregion, )) must be considered for accurate timing estimations. Planners must accurately model compute, memory, and network bandwidth heterogeneity - yet, as shown in Table 1, most systems overlook this.
@@
C3: Supporting both heterogeneous plans and seamless elasticity in a real distributed training framework. Most training frameworks, i.e., systems that train a model on a set of devices, assume homogeneous clusters and job configuration plans. For example, widely used and high-performing frameworks such as Megatron (shoeybi2020megatronlm, ) and DeepSpeed (megatron-deepspeed, ) assume uniform parallelism degrees for the entire training job (e.g. DP=2, PP=6, TP=1). This limits efficiency in heterogeneous configurations, where various parallelism degrees per training subproblem (e.g., using PP=6, with first 3 stages having TP=4, and the next 3 stages having TP=2) help load-balance compute and memory on GPU nodes with different resources. Thus, a framework should accommodate heterogeneous degrees of parallelism. Furthermore, the framework should seamlessly reconfigure and adapt to dynamic GPU availability (Figure 2), as long job reconfiguration times in response to resource changes are wasteful (Thorpe23Bamboo, ; wagenlander24tenplex, ). Although related works propose elastic systems (Athlur22Varuna, ; Thorpe23Bamboo, ; Duan24Parcae, ) or systems that support heterogeneity (Miao23SDPipe, ; yan2024flashflex, ), there is no open-source system that supports both.
@@ 4. Sailor
To address the above challenges, we propose Sailor, a distributed training ecosystem that consists of a profiler, planner, simulator, and distributed training framework. As shown in Figure 4, ML developers submit their model training specifications (model, optimizer, global batch size, etc), resource quotas (the maximum number of GPUs for each type and zone), an objective (e.g., maximize throughput or minimize cost), and optionally also constraints (e.g, a maximum budget per iteration or a minimum iteration time). Sailor also receives feedback about the current availability of hardware resources (which may be less than the quotas).
Workflow. The Sailor profiler 1 collects information about the training job, the compute nodes and network bandwidth (§4.1). The Sailor planner 2 uses this information to select a near-optimal resource allocation from the pool of available hardware and a job parallelization plan that optimizes the user’s objective within the provided constraints (§4.2). The planner uses the simulator 3 to accurately evaluate various candidate plans (§4.3) in terms of throughput, memory footprint, and cost. Sailor then launches the job with the selected configuration using its distributed training framework 44.4), which is implemented on top of Megatron-DeepSpeed (megatron-deepspeed, ). Sailor dynamically re-configures the job as resource availability changes.
@@ 4.1. Sailor Profiler
Training job profiling: When a user submits a training job, the Sailor profiler collects information about the job’s compute and memory requirements. Sailor profiles a training job on a single GPU node for each different GPU node type in the available resource pool. To minimize profiling time and enable single-node profiling, the profiler reduces repeated layers to a single instance (e.g., it uses one transformer layer for a given LLM). We use PyTorch hooks (torch_hooks, ) to collect information per layer: the time required for the forward pass, backward pass, and update phase with different microbatch sizes and tensor parallel degrees. We use CUDA Events for accurate GPU measurements (torch_cuda_events, ). We also track the number of parameters, output activation, and the memory required for intermediate stages per layer, using the PyTorch CUDA memory allocator (torch_cuda_mem_alloc, ). The profiling overhead is negligible (a couple of minutes).
Cluster profiling: Sailor also collects information about the network bandwidth between any pair of different machine types. Since network bandwidth depends on the message size, Sailor collects network bandwidth measurements (using PyTorch collectives with NCCL backend) by varying the message size and fitting a polynomial function to get a set of coefficients for any pair of node types.
@@ 4.2. Sailor Planner
The Sailor planner takes as input the training job and cluster profiling information from the profiler, a performance or cost objective, and optionally a constraint such as a budget. The planner selects a resource allocation and a parallelization plan to optimize the objective under the constraints. The parallelization plan defines the number of pipeline stages $P$ (which we refer to simply as stages), the data parallelism degree of each stage $D$, the $D$ pairs of $(GPU_{j},TP_{j},Zone_{j})$ for each stage, and the microbatch size $mbs$. The Sailor planner does not change the global batch size, thus it does not affect the job’s training dynamics.
@@
The planner iterates through different parallelism degrees and microbatch sizes (based on model characteristics and profiling information). For a given layer partitioning and microbatch size, it finds the tensor parallelism degree for each GPU type, based on memory constraints and scaling heuristics. Then, it iterates through all cloud region combinations and selects the data parallelism degrees to evaluate for a combination. For a fixed data parallelism degree, the planner invokes the $solve\_dp()$ function (Listing 1) that applies dynamic programming to determine the optimal stage configuration (§4.2.2). The planner considers the configuration valid only when it is within the user-specified constraint.
@@ 4.2.1. Pruning the large search space with heuristics
H1: Limit tensor parallelism within a node. Tensor parallelism performance is known to degrade when spanning multiple nodes (shoeybi2020megatronlm, ; strati2024dejavu, ). We restrict tensor parallelism to a single node and do not explore cross-node pairs (unlike Metis (Um24Metis, )). As a result, each tensor parallel replica of a stage only uses a single GPU type.
H2: Prune OOM configurations early. Since each stage replica performs tensor parallelism only among GPUs of the same type, we can easily compute the minimum tensor parallelism degree of each GPU type, for a given pipeline parallel stage and microbatch size. We exclude cases with tensor parallelism below this minimum. The minimum tensor parallelism for each stage is independent of the number of available GPUs per type, so we can reuse it when resource availability changes.
@@
H5: Keep data parallel communication within a single region, while spreading pipeline parallel communication across more than one region. As shown by earlier work (strati24crossregion, ; palak2024atlas, ), data parallelism performs poorly across regions due to the low network bandwidth. We constraint all data parallel pairs of a stage within a single region.
H6: Treat multiple zones within the same region as a single zone. Within a cloud region, the network bandwidth across zones is similar to the network bandwidth within a zone (strati24crossregion, ). Thus, to reduce the search space, we consolidate all zones in a region into a single zone and do the geo-distributed partitioning at a region granularity.
@@ 4.2.2. Selecting per-stage configurations with dynamic programming
Problem formulation and goal: Given a pipeline parallel degree $P$, a microbatch size $mbs$, a data parallel degree $D$, and tensor parallel degrees $tp_{ij}$ for GPU $j$ and stage $i$, we want to find, for each stage $i$, the $D$ replicas, where each replica is a tuple $(j,tp_{ij},zone_{k})$ that minimizes iteration time. Note that Sailor precomputes $tp_{ij}$ (Heuristic H2).
Why dynamic programming? Optimizing resource allocation per stage is challenging, as assigning resources to one stage impacts overall runtime and availability for other stages. In a heterogeneous, multi-zone setup, the number of possible resource combinations per stage explodes. Related works on homogeneous clusters used Integer Linear Programming (ILP) (zheng2022alpa, ; Jang2023Oobleck, ), exhaustive search  (Athlur22Varuna, ), or dynamic programming (li2022amp, ; Tarnawski21Piper, ). We adopt dynamic programming for its ability to decompose the problem into subproblems and reuse intermediate results.
@@
(1) $T_{iter}[l_{0}][P][R]=min_{r}\{T_{total}(T_{iter}[l_{0}+L][P-1][R-r],T_{i}(r))\}$
for a given pipeline parallel degree $P$, stage $i$, and available resources $R$. We give $r$ resources to stage $i$, and $R-r$ resources are available for the subsequent stages. In Sailor, $r$ represents a map of different GPU types and zones, and should be enough to get $D$ replicas of this stage. Based on heuristic H5, we keep each stage within a single region. $T_{total}$ is calculated by identifying the straggler between stage $i$ and the rest pipeline, the pipeline communication cost between stage $i$ and $i+1$, and the synchronization bottleneck between stage $i$ and the rest pipeline.
@@ 4.2.3. Adding a monetary cost constraint
$\sum_{0}^{P-1}{Ccomp_{i}}\cdot(\sum_{0}^{P-1}{t_{i}}+N_{batches}\cdot t_{straggler}+max_{0}^{P}(t_{sync_{i}}))+C_{comm}<=C$
where $\sum_{0}^{P-1}{t_{i}}$ stands for the pipeline warmup and cooldown phase, $N_{batches}$ is the number of microbatches processed per pipeline, and $t_{sync_{i}}$ is the time required for the synchronization of all replicas of stage $i$. Since large models usually train with large global batch sizes, the $N_{batches}\cdot t_{straggler}$ term usually determines the iteration time, we can rewrite as:
(2) $\sum_{0}^{P-1}{C_{{comp}_{i}}}\cdot(N_{batches}\cdot t_{straggler})+C_{comm}<=C$
From a dynamic programming perspective, assuming we are at pipeline stage $i$, which has a maximum budget limit $C_{cur}$, the cost constraint will be: $C_{i}+C_{rem}<=C_{cur}$, where $C_{rem}$ is the cost of the remaining stages, and $C_{i}$ is the cost of stage $i$. From Eq. 2, we have that $C_{i}=C_{comp_{i}}\cdot(N_{batches}\cdot t_{straggler})+C_{comm_{i}}$. In Listing 1, line 14, when exploring a resource combination $r$, we can easily find $C_{comp_{i}}$, and $C_{comm_{i}}$ for stage $i$. Since we do not know $t_{straggler}$, which is required to specify the remaining budget for the next stages, we use the approximation in lines 17-32: We begin assuming stage $i$ is the straggler ($t_{straggler}==t_{i}$), and compute the $C_{rem}$ for the next stages accordingly. We call the solve_dp function for the next stages giving $C_{rem}$ as the budget constraint (line 20). If a solution cannot be found, we proceed with the next resource combination for stage $i$ (line 22). If a solution is found, we check whether it is within the cost limit and keep the one with the maximum throughput (line 26). We then check the straggler of the found solution: if it is the same with the one we assumed, we break, and proceed to the next resource combination (lines 27-28). Otherwise, we adjust the budget with the new straggler (lines 31-32) and iterate again.

### v2

@@ 3.2. Challenges with Heterogeneous ML Clusters
C1: Quickly searching a vast configuration space. Considering heterogeneous and geo-distributed GPUs creates a vast and complex search space. The ML developer needs to decide how many GPUs to use and how to group them across VMs. This complexity increases further, when accounting for job parallelization plans within each allocation. Furthermore, the optimal allocation and partitioning strategies depend on user objectives and constraints. A planner needs to quickly navigate the large configuration space to adapt to dynamic resource availability in cloud and on-premise environments (Thorpe et al., 2023; Guo et al., 2024; Weng et al., 2022; Gu et al., 2019), since maximizing throughput requires adapting parallelization strategies with changes to cluster topology (Athlur et al., 2022; Wagenländer et al., 2024).
Table 1 shows that Metis (Um et al., 2024), FlashFlex (Yan et al., 2024), Cephalo (Guo et al., 2024), Atlas (Palak et al., 2024a), and DTFM (Yuan et al., 2023) explore parts of this large search space. Atlas (Palak et al., 2024a) and DTFM (Yuan et al., 2023) target geo-distributed training, but do not consider heterogeneous GPU types, and do not decide the various parallelism degrees: instead, they take as input the parallelism degrees, and assign these degrees in the available zones. On the other hand, Metis (Um et al., 2024), FlashFlex (Yan et al., 2024) and Cephalo (Guo et al., 2024), consider heterogeneous GPU types, but overlook geo-distributed training, and are quite inefficient for dynamic environments. Metis needs a few hours to devise a plan for a 16-GPU cluster (A100 and V100)333with max_permutation_length and device group variance set to 10 and 0.5 respectively, according to the paper (Um et al., 2024), making frequent reevaluation infeasible as GPU availability changes. Cephalo (Guo et al., 2024) has a search time of  300 sec on a cluster of 64 GPUs444reported on the paper (Guo et al., 2024), but it is limited only to Fully Sharded Data Parallelism. FlashFlex (Yan et al., 2024) has a short runtime, but provides inaccurate runtime estimations, leading to suboptimal plans. Furthermore, these planners only optimize for throughput, ignoring budget constraints, and cost (dollars per iteration), which affect the optimal configuration (§5.2.4).
@@
Figure 3. Peak Memory estimations of various baselines compared to the actual peak memory, for the OPT-350M model on a homogeneous cluster of 4 Grace-Hopper per node. NN stands for the number of nodes, gbsgbs is the global batch size, pppp pipeline parallelism, tptp tensor parallelism, dpdp data parallelism, and mbsmbs the microbatch size.
@@
Unfortunately, these estimations are often inaccurate, resulting in suboptimal or invalid plans. First, planners either completely ignore memory footprint (Li et al., 2022), or underestimate the amount of memory requirements during training (Athlur et al., 2022; Jang et al., 2023), omitting activations, optimizer states, and memory fragmentation, or assume the training memory footprint is uniform across all devices and pipeline stages (Yan et al., 2024; Lin et al., 2024). As a result, these systems may find plans that cause OOM errors, when deployed. Figure 3 compares memory footprint estimations with the real footprint on a homogeneous cluster of up to 16 Grace-Hopper nodes for the OPT-350M model, showing that planners can be 25-95% off when estimating memory footprint. Second, planners often poorly model training time, due to incorrect assumptions about network bandwidth and ignoring communication-computation overlap (Zhang et al., 2017; deepspeedai, 2025).
@@
Modeling memory footprint and iteration time becomes even more complex with heterogeneity (Mo et al., 2024). Different GPU generations vary in compute performance (e.g., TFLOPs), and memory capacity (Zhu et al., 2024). Thus, a configuration that fits in one GPU, might cause OOM errors in another GPU. Additionally, stragglers and network bandwidth differences (especially in geo-distributed setups (Strati et al., 2024a)) must be considered for accurate timing estimations. Planners must accurately model compute, memory, and network bandwidth heterogeneity - yet, as shown in Table 1, most systems overlook this.
@@
C3: Supporting both heterogeneous plans and seamless elasticity in a real distributed training framework. Most training frameworks, i.e., systems that train a model on a set of devices, assume homogeneous clusters and job configuration plans. For example, widely used and high-performing frameworks such as Megatron (Shoeybi et al., 2020) and DeepSpeed (deepspeedai, 2025) assume uniform parallelism degrees for the entire training job (e.g. DP=2, PP=6, TP=1). This limits efficiency in heterogeneous configurations, where various parallelism degrees per training subproblem (e.g., using PP=6, with first 3 stages having TP=4, and the next 3 stages having TP=2) help load-balance compute and memory on GPU nodes with different resources. Thus, a framework should accommodate heterogeneous degrees of parallelism. Furthermore, the framework should seamlessly reconfigure and adapt to dynamic GPU availability (Figure 2), as long job reconfiguration times in response to resource changes are wasteful (Thorpe et al., 2023; Wagenländer et al., 2024). Although related works propose elastic systems (Athlur et al., 2022; Thorpe et al., 2023; Duan et al., 2024) or systems that support heterogeneity (Miao et al., 2023; Yan et al., 2024), there is no open-source system that supports both.
@@ 4. Sailor
To address the above challenges, we propose Sailor, a distributed training ecosystem that consists of a profiler, planner, simulator, and distributed training framework. As shown in Figure 4, ML developers submit their model training specifications (model, optimizer, global batch size, etc), resource quotas (the maximum number of GPUs for each type and zone), an objective (e.g., maximize throughput or minimize cost), and optionally also constraints (e.g, a maximum budget per iteration or a maximum iteration time). Sailor also receives feedback about the current availability of hardware resources (which may be less than the quotas).
Workflow. The Sailor profiler 1 collects information about the training job, the compute nodes and network bandwidth (§4.1). The Sailor planner 2 uses this information to select a near-optimal resource allocation from the pool of available hardware and a job parallelization plan that optimizes the user’s objective within the provided constraints (§4.2). The planner uses the simulator 3 to accurately evaluate various candidate plans (§4.3) in terms of throughput, memory footprint, and cost. Sailor then launches the job with the selected configuration using its distributed training framework 44.4), which is implemented on top of Megatron-DeepSpeed (deepspeedai, 2025). Sailor dynamically re-configures the job as resource availability changes.
@@ 4.1. Sailor Profiler
Training job profiling: When a user submits a training job, the Sailor profiler collects information about the job’s compute and memory requirements. Sailor profiles a training job on a single GPU node for each different GPU node type in the available resource pool. To minimize profiling time and enable single-node profiling, the profiler reduces repeated layers to a single instance (e.g., it uses one transformer layer for a given LLM). We use PyTorch hooks (PyTorch, 2025b) to collect information per layer: the time required for the forward pass, backward pass, and update phase with different microbatch sizes and tensor parallel degrees. We use CUDA Events for accurate GPU measurements (PyTorch, 2025a). We also track the number of parameters, output activation, and the memory required for intermediate stages per layer, using the PyTorch CUDA memory allocator (PyTorch, 2025c). The profiling overhead is negligible (a couple of minutes) for the LLMs we consider. For models with non-uniform layers, all layers need to be profiled, which increases profiling time. Our profiling approach can be used with any dense layers, while profiling Mixture-of-Experts (where layer load may vary dynamically during training) is left for future work.
Cluster profiling: Sailor also collects information about the network bandwidth between any pair of different machine types. Since network bandwidth depends on the message size, Sailor collects network bandwidth measurements (using PyTorch collectives with NCCL backend) by varying the message size and fitting a polynomial function to get a set of coefficients for any pair of node types. Our profiling methodology applies to all GPU types. Adding a new GPU type requires collecting model and cluster profiling data as described above.
@@ 4.2. Sailor Planner
The Sailor planner takes as input the training job and cluster profiling information from the profiler, a performance or cost objective, and optionally a constraint such as a budget. The planner selects a resource allocation and a parallelization plan to optimize the objective under the constraints. The parallelization plan defines the number of pipeline stages $P$ (which we refer to simply as stages), the data parallelism degree of each stage $D$, the $D$ pairs of $(GPU_{j},TP_{j},Zone_{j})$ for each stage, and the microbatch size mbsmbs. The Sailor planner does not change the global batch size, thus it does not affect the job’s training dynamics.
@@
The planner iterates through different parallelism degrees and microbatch sizes (based on model characteristics and profiling information). For a given layer partitioning and microbatch size, it finds the tensor parallelism degree for each GPU type, based on memory constraints and scaling heuristics. Then, it iterates through all cloud region combinations and selects the data parallelism degrees to evaluate for a combination. For a fixed data parallelism degree, the planner invokes the solve_dp()solve\_dp() function (Listing 1) that applies dynamic programming to determine the optimal stage configuration (§4.2.2). The planner considers the configuration valid only when it is within the user-specified constraint.
@@ 4.2.1. Pruning the large search space with heuristics
H1: Limit tensor parallelism within a node. Tensor parallelism performance is known to degrade when spanning multiple nodes (Shoeybi et al., 2020; Strati et al., 2024b). We restrict tensor parallelism to a single node and do not explore cross-node pairs (unlike Metis (Um et al., 2024)). As a result, each tensor parallel replica of a stage only uses a single GPU type.
H2: Prune OOM configurations early. Since each stage replica performs tensor parallelism only among GPUs of the same type, we can easily compute the minimum tensor parallelism degree of each GPU type, for a given pipeline parallel stage and microbatch size. We exclude cases with tensor parallelism below this minimum. To find the minimum tensor parallelism degree per GPU for a stage, we compute the memory footprint of that stage as described in §4.3, and identify the minimum number of GPUs required based on the available memory per GPU. The minimum tensor parallelism for each stage is independent of the number of available GPUs per type, so we can reuse it when resource availability changes.
@@
H5: Keep data parallel communication within a single region, while spreading pipeline parallel communication across more than one region. As shown by earlier work (Strati et al., 2024a; Palak et al., 2024a), data parallelism performs poorly across regions due to the low network bandwidth. We constraint all data parallel pairs of a stage within a single region.
H6: Treat multiple zones within the same region as a single zone. Within a cloud region, the network bandwidth across zones is similar to the network bandwidth within a zone (Strati et al., 2024a). Thus, to reduce the search space, we consolidate all zones in a region into a single zone and do the geo-distributed partitioning at a region granularity.
@@ 4.2.2. Selecting per-stage configurations with dynamic programming
Problem formulation and goal: Given a pipeline parallel degree $P$, a microbatch size mbsmbs, a data parallel degree $D$, and tensor parallel degrees $tp_{ij}$ for GPU $j$ and stage $i$, we want to find, for each stage $i$, the $D$ replicas, where each replica is a tuple $(j,tp_{ij},zone_{k})$ that minimizes iteration time. Note that Sailor precomputes $tp_{ij}$ (Heuristic H2).
Why dynamic programming? Optimizing resource allocation per stage is challenging, as assigning resources to one stage impacts overall runtime and availability for other stages. In a heterogeneous, multi-zone setup, the number of possible resource combinations per stage explodes. Related works on homogeneous clusters used Integer Linear Programming (ILP) (Zheng et al., 2022; Jang et al., 2023), exhaustive search  (Athlur et al., 2022), or dynamic programming (Li et al., 2022; Tarnawski et al., 2021). We adopt dynamic programming for its ability to decompose the problem into subproblems and reuse intermediate results.
@@
[…]
for a given pipeline parallel degree $P$, stage $i$, and available resources $R$. We give $r$ resources to stage $i$, and RrR-r resources are available for the subsequent stages. In Sailor, $r$ represents a map of different GPU types and zones, and should be enough to get $D$ replicas of this stage. Based on heuristic H5, we keep each stage within a single region. $T_{total}$ is calculated by identifying the straggler between stage $i$ and the rest pipeline, the pipeline communication cost between stage $i$ and $i+1$, and the synchronization bottleneck between stage $i$ and the rest pipeline.
More specifically, assuming XiX-i is the pipeline without stage $i$, and $N_{b}$ is the number of microbatches processed per pipeline, $t_{j}$ is the time per stage, and $sync_{j}$ is the time for synchronization:
$T_{total}=max(t_{i}(r),Straggler_{X-i}(R-r))\cdot N_{b}\newline$
[…]
Our formulation aims to minimize iteration time and reduce pipeline stragglers, inherently excluding imbalanced pipelines with large stage time differences.
@@ 4.2.3. Adding a monetary cost constraint
[…]
From a dynamic programming perspective, assuming we are at pipeline stage $i$, which has a maximum budget limit $C_{cur}$, the cost constraint will be: $C_{i}+C_{rem}<=C_{cur}$, where $C_{rem}$ is the cost of the remaining stages, and $C_{i}$ is the cost of stage $i$. From Eq. 2, we have that $C_{i}=C_{comp_{i}}\cdot(N_{b}\cdot t_{straggler})+C_{comm_{i}}$. In Listing 1, line 14, when exploring a resource combination $r$, we can easily find $C_{comp_{i}}$, and $C_{comm_{i}}$ for stage $i$. Since we do not know $t_{straggler}$, which is required to specify the remaining budget for the next stages, we use the approximation in lines 17-32: We begin assuming stage $i$ is the straggler ($t_{straggler}==t_{i}$), and compute the $C_{rem}$ for the next stages accordingly. We call the solve_dp function for the next stages giving $C_{rem}$ as the budget constraint (line 20). If a solution cannot be found, we proceed with the next resource combination for stage $i$ (line 22). If a solution is found, we check whether it is within the cost limit and keep the one with the maximum throughput (line 26). We also check the straggler of the found solution: if it is the same with the one we assumed, we break, and proceed to the next resource combination (lines 26-28). Otherwise, we adjust the budget with the new straggler (lines 31-32) and iterate again.